Efficient parallel solving of the potential flow problem on nonmatching grids

Yu.Iliash  Yu.Kuznetsov  Yu.Vassilevski

Institute of Numerical Mathematics RAS, Leninskij prosp., 32a, Moscow 117334, Russia.

Abstract

We consider a parallel application of DD algorithm proposed in [1] to approximate solving of fully potential flow problem on nonmatching grids. A short description of the method and results of numerical experiments on parallel computer SP2 are presented.

References

[1] Yu. Kuznetsov, Efficient iterative solvers for elliptic finite element problems on nonmatching grids, Russ.J.Numer.Anal.Math.Modelling, 10, No.3, 187–211 (1995).